HETAV PATEL

Grade 11,

Consider an ice cube of edge 1.0 cm kept in a gravity free hall. Find the surface area of the water when the ice melts. Neglect the difference in densities of ice and water.

Consider an ice cube of edge 1.0 cm kept in a gravity free hall. Find the surface area of the water when the ice melts. Neglect the difference in densities of ice and water.

Grade:10

4 Answers

Deepak Patra
9 years ago
Sol. Surface area = 4πr2
Koustav Saha
9 years ago
Consider an ice cube of edge 1.0 cm kept in a gravity free hall.Find the surface area of water when the ice melts
 
Subodh Wasnik
6 years ago
I am unable to solve it. I tried very but cannot get the steps for it. However i know the answer. Answer is (9π)^1/3
Subodh Wasnik
6 years ago
Since ice is in gravity free space, thus after melting it forms a sphere. Area of ice cube= area of sphere1 cm3 =4/3πr^3=> r=(3/4π)^1/3Now, surfacr area is given byS.A.= 4πr^2 =4π(3/4π)^2/3 =(9π)^1/3